y=-0.0575x+7.3


1
Expert's answer
2021-09-28T12:58:51-0400

Regression equation y=-0.0575x+7.3 is of the form y=a*x+ b, where a= -0.0575 is the slope coefficient and b=7.3 is the y-intercept.

The slope coefficient of -0.0575 represents the estimated decrease in y when x increases by one unit.

The y-intercept= 7.3 is the expected mean value of y when all x=0.